• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jin Huang

    Log entry time 23:49:49 on October 7,2010

    Entry number 333904

    keyword=Target Screen 6

    Hall C 15K flow has been dropping. Seem done by ESR. But both supply and return temperature still remain high. Target loop temperature has also been stable.

    Current status:
    Loop1: 28K
    Loop2: 71K
    15K sup.: 18K
    20K ret.: 33K

    My suggest to Owl target operator to watch:
    1. 20K return temperature (CTD674): stable and <~40K
    2. Loop 1 temperature (haL1_Cernox_1/2/3_R): stable. Could close loop1 JT valve a little if it drop <19K.



    FIGURE 1

    FIGURE 2